# Analysis Theory of Data Economy : Dataization, Technological Progress and Dynamic General Equilibrium

### v1

@@ Abstract
This paper constructs a clean and efficient representative agent model of the data economy from a macroeconomics perspective, in order to analyze the impact of datatization and technological progress on the dynamic general equilibrium of “consumption-capital”, and the catalysis effect of datatization on technological progress. We first set the data in production comes from datatization of the total output of the society and is exponentially functionally related to the technology. Secondly, the data production function is used to solve the optimization problem for firms and households and to construct a dynamic general equilibrium of “consumption-capital” based on the endogenous interest rate solved by maximizing the returns of firms. Finally, by using numerical simulation and phase diagram analysis, we find that the effects of increasing datatization and encouraging technological progress each exhibit different nonlinear characteristics for equilibrium capital and equilibrium consumption, we thus conclude that datatization enables the positive effects of technological progress on economic development to be more rapid and persistent. We select two types of Chinese policies regarding data openness to represent the role of datatization, and demonstrate the catalysis role of datatization for the development of the digital economy by setting up difference in difference (DID) experiments, which provide persuasive evidence for the theoretical interpretation of the paper.
Key Words: Data economy, Datatization, Technological progress, Representative agent model, Dynamic general equilibrium
@@ 1 Introduction
Data essentially shifts the material basis of knowledge and information production from traditional discourse practices to quantifiable and tradable data resources, completing the datatization transformation of knowledge, information and production factors. The special economic attributes of data as a factor of production, such as non-rivalry, replicability, virtuality, open source, and non-depletion at the physical level (Jones and Tonetti, 2020; Goldfarb and Tucker, 2019; Lynch, 2008), enable it to be integrated into the cycle of capital accumulation in a way of decreasing marginal cost and increasing returns to scale.
@@
Data collection improves services and benefits users, but users also face security issues such as privacy breaches (Cong and Mayer, 2023), which in turn incur privacy costs derived from data oversharing (Cong et al., 2021; Jones and Tonetti, 2020). With regard to the use and regulation of big data, e-commerce online platforms may use consumer information materialized by data elements to alter the “gloss” of product quality, causing consumers to misjudge the true quality of the product and inducing unwanted consumption behaviors (Acemoglu et al., 2025), Moreover, platforms will not only use the pricing mechanism of “big data to kill maturity” to make profits, but may also redistribute information in the form of “filter bubbles” for heterogeneous personal data in order to maximize platform engagement (Acemoglu et al., 2024). A study by Ding et al. (2024) argues that the stimulus effect on consumption growth of consumer benefits translated from data elements, such as the issuance of digital consumption vouchers or shopping subsidies, stems from the increase in consumer spending in the targeted category and not from crowding out of other categories of consumption expenditure. Madsen and Vellodi (2025) point out that a complete ban on data use may foster or inhibit innovation, depending on the demand for new products, and therefore more flexible policies that control when and what data are made available are needed to increase the effectiveness of data regulation. Some studies have incorporated information uncertainty into the “efficiency-equity” research framework, pointing out that information frictions connect data generation and economic activity, mainly in the following ways: micro-level uncertainty creates resource mismatches in the macrosystem, making macro total factor productivity endogenous to micro agents’ data collection behavior (Farboodi and Veldkamp, 2021; David et al., 2016; Benhabib et al., 2016). This may be due to the widespread use of facilitated mobile communication tools and the availability of data information that allows for more random and rapid changes in individual behavior (Fabregas et al., 2025), which leads to the inability of traditional data selection mechanisms to accurately identify imperfect information and decision-making errors, as Gans (2025b) points out that in domains rich in datatization variety, AI’s analytical and decision-making capabilities performs well, while it is less credible in judgment-intensive data environments. Therefore, Caplin (2025) introduces new forms of data to identify agents’ preferences, beliefs, etc. by constructing data engineering models. Notably, Jones (2025) and Jones (2024) explore the potential for big data-enabled AI technologies to pose a threat to human existence while promoting economic growth at the level of heterogeneous agent utility acquisition.
@@
To summarize, there are two mainstream conclusive studies in academia: the first one views data as a form of medium for information dissemination, which determines the source and value of information (Farboodi and Veldkamp, 2021); and the second one views data as an input factor, which determines technological advancement and output (Jones and Tonetti, 2020). And our data economy [^2] The framework of our paper is based on the perspective of “data economy”, which is different from digital economy. Data economy is a form of economic research centered on data as a factor of production, which centers on value creation through the production, distribution, exchange and consumption of data resources. Data economy emphasizes that data, as a non-rivalry, reusable resource, can be embedded in the production function to optimize the efficiency of resource allocation and enhance productivity. Its theoretical roots can be traced back to the breakthrough of the assumption of technological exogeneity in neoclassical growth theory and the extension of knowledge spillovers in Romer’s (1990) endogenous growth theory. Digital economy refers to the new economic system reconstructed on the basis of Digital Technologies, covering the two dimensions of digital infrastructure, digital industrialization and industrial digitization. Its substance is the use of digital general-purpose technology, triggered by the total factor productivity increase and organizational model change, manifested in the expansion of the production boundary and the reduction of transaction costs. analysis refers to Jones and Tonetti (2020), viewing data as a factor of production to study its impact on the economy.
@@
Marginal contributions to the article: First, we analyze the impact of datatization and technological progress on the dynamic general equilibrium of “consumption-capital” and summarize the conclusions about the development of data economy through numerical simulation and phase diagram analysis, filling the gap in the basic theory of this field. Secondly, most of the existing articles focus on the impact of datatization shock or the proliferation of data technologies on the economy, ignoring the role of datatization as a catalyst for technological progress from a theoretical perspective. Therefore, the article specifically explores the relationship between datamization and technological progress and argues for it using empirical analyses, providing recommendations for achieving faster, more persistent and more effective technological progress in the data economy.
@@ 2.1 The Production Function of Data
The Prerequisite. We assume that all consumers and firms in the macroeconomy are homogeneous, i.e., representative consumers and representative firms. In this macroeconomy, the consumption of representative consumers and the investment of representative firms will generate big data, and these data elements can be collected and applied to the study of consumer consumption behavior and firm investment behavior, and to the formation of algorithms and techniques. The total output of the macroeconomy comes from the economic information generated by the decisions of representative consumers and firms, and as to what proportion of the total output can be legally datatized and applied to research, this paper sets out to show that this is an exogenous variable that varies with changes in digital economy policies, which will be determined by the policies set by the government. That is, policies that do not allow some of the data to be applied to algorithms for studying consumer and business behavior due to the dual factors of enhancing consumer welfare and protecting individual privacy, or, policies that stipulate that only no more than a certain percentage of the total output can be applied to the development of technology and innovation.
@@
Now, taking data as an element of the production function, based on Jones and Tonetti (2020), assume an exponential relationship $\eta$ between data $d(t)$ and technology $z(t)$. Introducing the exogenous variable $\theta$, assume that the policy stipulates that only $\theta$ proportions of output can be datatized and put into production as data elements, i.e.:
@@
It is important to note that the exponential transformation rate relationship between data and technology stems from the nonrivalry and cumulative effects of data. As a factor of production, data can be shared indefinitely without loss, and its value grows exponentially with the scope of its use. In the technical production function, the amount of data directly affects the production efficiency of technology through the exponential form $z(t)=d(t)^{\eta},\eta=\left(\frac{logz(t)}{logd(t)}\right)$, reflecting the importance of data to technology. As the amount of data increases, the output efficiency of technology increases at an exponential rate, forming a positive cycle: more data→higher technical efficiency→more output→generate more data.For example, the training of autonomous driving techniques requires a large amount of driving data. Assuming that an algorithm uses 1 million kilometers of driving data with 80% accuracy, when the amount of data is increased to 10 million kilometers (10 times), the accuracy may be increased to 95% due to the cumulative effect of non-competitive data (non-linear growth). This improvement is not linearly cumulative, but rather the data is trained by the algorithm to form an exponential optimization, reflecting the exponential transformation relationship between data and technology.
@@ 3 Numerical Simulation and Phase Diagram Analysis
From Figure 1-A: The effect of changes in $\eta$ and $\theta$ on equilibrium investment $k^{*}$ is linear. An increase in the data technology transformation rate $\eta$ promotes investment, while an increase in the proportion of total output datatization $\theta$ depresses investment. Thus, the boosting effect of $\eta$ on the equilibrium capital stock will gradually slow down with an increase in $\theta$. Figure 1-B shows that the impact of changes in the data technology conversion rate $\eta$ and the proportion of total output datatization $\theta$ on the equilibrium consumption $c^{*}$ is non-linear and has a “mixing” effect. $c^{*}$ is affected by $\eta$ at an inflection point, and the inflection point moves backward with increasing $\theta$. This implies that the promotion of consumption by $\eta$ is not persistent, and there is a possibility that the equilibrium consumption level will decrease at higher stages of data technology conversion. We interpret this as follows:
@@
On the one hand, it is clear from Figure 1-A that technological progress can increase the marginal productivity of capital factors thereby driving the equilibrium capital stock, which is consistent with the rules of reality. And when the datatization ratio of total output increases, from an economic perspective, higher $\theta$ implies that more data are produced, and technology significantly improves and raises total factor productivity, at this time, firms are more biased toward data-driven rather than capital factor inputs for technological improvements, and technological improvement contributes to output partially substituting for capital, coupled with a shift in resources from physical capital investment to data production, leading to a decrease in capital required in the steady state. And, in an environment where technology is highly sensitive to data, i.e., the higher the level of $\eta$, the more pronounced is the impact of the equilibrium capital stock by the increase in the proportion of datatization. As can be seen from Figure 1-B, when $\eta$ is small, an increase in $\theta$ does not have a significant effect on equilibrium consumption. In contrast, when $\eta$ is large, an increase in $\theta$ has the effect of promoting equilibrium consumption. Since $\eta$ is the data-technology conversion rate, which represents the maturity of the mechanism for converting data into technology, the larger $\eta$ is, the more mature the conversion mechanism is, the more the same amount of data will be converted into a higher level of technological achievements, thus more efficiently enabling the transmission of initial data from the production side to the demand side of the individual. Consumers also utilize data-forming technologies to assist in decision-making and effective consumption. Therefore, with a mature mechanism of data technology conversion, increasing the proportion of total macroeconomic output datatization can stimulate consumption growth.
@@
On the other hand, Figure 1-B shows that the overall impact of $\eta$ on equilibrium consumption shows an inverted “U”-shaped trend, which implies that transforming big data into technology is not always effective in stimulating consumption growth, i.e., there exists a threshold for the impact of an increase in the data-technology conversion rate on equilibrium consumption, and the size of this threshold is related to The size of this threshold is related to the proportion of total output datatization $\theta$. Specifically, when $\theta$ is larger, in the domain of definition of $\eta$, the range of promoting effect of increasing $\eta$ on $c^{*}$ will be more persistent, and the threshold of the inverted “U”-shaped trend will be more backward, and gradually evolve into “S”-shaped logistic curve. This may be due to the scale of data is the basis for the establishment of the data-technology transformation mechanism (Learning from Big Data). Therefore, the proportion of total output datatization increases, which means that after the total output is datatized, the volume of data that can be used for technological transformation also increases, and the space and resources for “learning” from data will be expanded, the more reliable the data-technology transformation mechanism is, the higher the success rate of data transformation into technology, and the more accurate the prediction of technology for consumer behavior, thus driving the “resilience” of effective consumption growth.
@@
Further examining Figure 1, we also find that both for $k^{*}$ and $c^{*}$, when $\eta$ is at a higher level ($0.6<\eta<1)$, there is a positive correlation between $\theta$ and $\eta$ at the same level of equilibrium (The Same Color Area in the Figure 1). Therefore, when consumption or capital stock is at the same level of steady state, increasing the macroeconomic total output datatization ratio $\theta$ can effectively promote technical progress $\eta$. Specifically as shown in Figure 2. This will be the theoretical basis for our empirical analysis.
@@
Phase Diagram Analysis. Figures 3, 4 further present the dynamic general equilibrium subject to shocks from changes in $\eta$ and $\theta$ in the macroeconomy of this paper, using the phase diagram approach:
@@
Figure 3 presents the impact of an increase in the data-technology conversion rate $\eta$ on consumption and capital stock in general dynamic equilibrium for a constant datatization ratio $\theta$ of total macroeconomic output: The black curve represents the case before the positive shock to $\eta$, $E_{1}$ represents the equilibrium point in that case, and $P_{1}$ represents the inflection point in that case; The blue curve represents the case at the beginning of the positive shock to $\eta$, with a small increase in $\eta$, $E_{2}$ represents the equilibrium point in that case, and $P_{2}$ represents the inflection point in that case; The red curve represents the case at the middle to late stage of the positive shock to $\eta$, with a larger increase in $\eta$, $E_{3}$ represents the equilibrium point, and $P_{3}$ represents the inflection point in that case; The three inflection points $P_{1}$,$P_{2}$,$P_{3}$ are covariant since the aggregate output datatization ratio $\theta$ is constant. When $\eta$ increases, equilibrium capital stock will increase linearly, and the phase diagram is shifted to the right. The blue phase diagram indicates that an increase in $\eta$ before the inflection point value will drive equilibrium consumption growth, and the red phase diagram indicates that an increase in $\eta$ after the inflection point value will slightly reduce equilibrium consumption.
@@
As can be seen from Figures 1-B, the impact of increasing datatization ratio $\theta$ on equilibrium consumption is not apparent in the stage of lower $\eta$. Therefore, the impact of increasing macroeconomic total output datatization ratio $\theta$ on consumption and capital stock in the dynamic general equilibrium in Figure 4 is analyzed for the stage of higher $\eta$: The black curve represents the case before the positive shock of theta, $E_{1}$ represents the equilibrium point in that case, and $P_{1}$ represents the inflection point in that case; The red curve represents the case when $\eta$ is higher as $\theta$ increases, $E_{2}$ represents the equilibrium point in that case, and $P_{2}$ represents the inflection point in that case; The two inflection points, $P_{1}$,$P_{2}$, increase sequentially to the right as the datatization ratio of total output $\theta$ continues to increase. As $\theta$ increases, the equilibrium capital stock will be linearly lower, and the phase diagram is shifted left. The red phase diagram indicates that expanding the proportion of datatization at the mature stage of data-technology conversion will be able to effectively stimulate consumption growth.
From figure 3 we can know that when the datatization ratio $\theta$ is fixed, it means that the threshold of $\eta$ to promote $c^{*}$ growth is fixed, and at this time a small increase in $\eta$ will effectively promote $c^{*}$ growth, while once the threshold is exceeded, continuing to raise $\eta$ will not be conducive to $c^{*}$ growth. And regardless of the magnitude of $\eta$ raising, its effect on $k^{*}$ is always promoting. From Figure 4 we can know that when eta is at a higher stage, raising the proportion of datatization $\theta$ will be able to effectively promote the $c^{*}$ growth and reduce $k^{*}$, at the same time, the threshold value of the effect of $\eta$ raising on the growth of $c^{*}$ will also increase because of $\theta$ increasing.

### v2

@@ Abstract
This paper constructs a clean and efficient representative agent model of the data economy from a macroeconomics perspective, in order to analyze the impact of dataization and technological progress on the dynamic general equilibrium of “consumption-capital”, and the catalysis effect of dataization on technological progress. We first set the data in production comes from dataization of the total output of the society and is exponentially functionally related to the technology. Secondly, the data production function is used to solve the optimization problem for firms and households and to construct a dynamic general equilibrium of “consumption-capital” based on the endogenous interest rate solved by maximizing the returns of firms. Finally, by using numerical simulation and phase diagram analysis, we find that the effects of increasing dataization and encouraging technological progress each exhibit different nonlinear characteristics for equilibrium capital and equilibrium consumption, we thus conclude that dataization enables the positive effects of technological progress on economic development to be more rapid and persistent. We select two types of Chinese policies regarding data openness to represent the role of dataization, and demonstrate the catalysis role of dataization for the development of the digital economy by setting up difference in difference (DID) experiments, which provide persuasive evidence for the theoretical interpretation of the paper.
Key Words: Data economy, Dataization, Technological progress, Representative agent model, Dynamic general equilibrium
@@ 1 Introduction
Data essentially shifts the material basis of knowledge and information production from traditional discourse practices to quantifiable and tradable data resources, completing the dataization transformation of knowledge, information and production factors. The special economic attributes of data as a factor of production, such as non-rivalry, replicability, virtuality, open source, and non-depletion at the physical level (Jones and Tonetti, 2020; Goldfarb and Tucker, 2019; Lynch, 2008), enable it to be integrated into the cycle of capital accumulation in a way of decreasing marginal cost and increasing returns to scale.
@@
Data collection improves services and benefits users, but users also face security issues such as privacy breaches (Cong and Mayer, 2023), which in turn incur privacy costs derived from data oversharing (Cong et al., 2021; Jones and Tonetti, 2020). With regard to the use and regulation of big data, e-commerce online platforms may use consumer information materialized by data elements to alter the “gloss” of product quality, causing consumers to misjudge the true quality of the product and inducing unwanted consumption behaviors (Acemoglu et al., 2025), Moreover, platforms will not only use the pricing mechanism of “big data to kill maturity” to make profits, but may also redistribute information in the form of “filter bubbles” for heterogeneous personal data in order to maximize platform engagement (Acemoglu et al., 2024). A study by Ding et al. (2024) argues that the stimulus effect on consumption growth of consumer benefits translated from data elements, such as the issuance of digital consumption vouchers or shopping subsidies, stems from the increase in consumer spending in the targeted category and not from crowding out of other categories of consumption expenditure. Madsen and Vellodi (2025) point out that a complete ban on data use may foster or inhibit innovation, depending on the demand for new products, and therefore more flexible policies that control when and what data are made available are needed to increase the effectiveness of data regulation. Some studies have incorporated information uncertainty into the “efficiency-equity” research framework, pointing out that information frictions connect data generation and economic activity, mainly in the following ways: micro-level uncertainty creates resource mismatches in the macro-system, making macro total factor productivity endogenous to micro agents’ data collection behavior (Farboodi and Veldkamp, 2021; David et al., 2016; Benhabib et al., 2016). This may be due to the widespread use of facilitated mobile communication tools and the availability of data information that allows for more random and rapid changes in individual behavior (Fabregas et al., 2025), which leads to the inability of traditional data selection mechanisms to accurately identify imperfect information and decision-making errors, as Gans (2025b) points out that in domains rich in dataization variety, AI’s analytical and decision-making capabilities performs well, while it is less credible in judgment-intensive data environments. Therefore, Caplin (2025) introduces new forms of data to identify agents’ preferences, beliefs, etc. by constructing data engineering models. Notably, Jones (2025) and Jones (2024) explore the potential for big data-enabled AI technologies to pose a threat to human existence while promoting economic growth at the level of heterogeneous agent utility acquisition.
@@
To summarize, there are two mainstream conclusive studies in academia: the first one views data as a form of medium for information dissemination, which determines the source and value of information (Farboodi and Veldkamp, 2021); and the second one views data as an input factor, which determines technological advancement and output (Jones and Tonetti, 2020). And our data economy [^2] The framework of our paper is based on the perspective of “data economy”, which is different from digital economy. Data economy is a form of economic research centered on data as a factor of production, which centers on value creation through the production, distribution, exchange and consumption of data resources. Data economy emphasizes that data, as a non-rivalry, reusable resource, can be embedded in the production function to optimize the efficiency of resource allocation and enhance productivity. Its theoretical roots can be traced back to the breakthrough of the assumption of exogenous technology in neoclassical growth theory and the extension of knowledge spillovers in Rome’s (1990) endogenous growth theory. Digital economy refers to the new economic system reconstructed on the basis of Digital Technologies, covering the two dimensions of digital infrastructure, digital industrialization and industrial digitization. Its substance is the use of digital general-purpose technology, triggered by the total factor productivity increase and organizational model change, manifested in the expansion of the production boundary and the reduction of transaction costs. analysis refers to Jones and Tonetti (2020), viewing data as a factor of production to study its impact on the economy.
@@
Marginal contributions to the article: First, we analyze the impact of dataization and technological progress on the dynamic general equilibrium of “consumption-capital” and summarize the conclusions about the development of data economy through numerical simulation and phase diagram analysis, filling the gap in the basic theory of this field. Secondly, most of the existing articles focus on the impact of dataization shock or the proliferation of data technologies on the economy, ignoring the role of dataization as a catalyst for technological progress from a theoretical perspective. Therefore, the article specifically explores the relationship between dataization and technological progress and argues for it using empirical analyses, providing recommendations for achieving faster, more persistent and more effective technological progress in the data economy.
@@ 2.1 The Production Function of Data
The Prerequisite. We assume that all consumers and firms in the macro-economy are homogeneous, i.e., representative consumers and representative firms. In this macro-economy, the consumption of representative consumers and the investment of representative firms will generate big data, and these data elements can be collected and applied to the study of consumer consumption behavior and firm investment behavior, and to the formation of algorithms and techniques. The total output of the macro-economy comes from the economic information generated by the decisions of representative consumers and firms, and as to what proportion of the total output can be legally dataized and applied to research, this paper sets out to show that this is an exogenous variable that varies with changes in digital economy policies, which will be determined by the policies set by the government. That is, policies that do not allow some of the data to be applied to algorithms for studying consumer and business behavior due to the dual factors of enhancing consumer welfare and protecting individual privacy, or, policies that stipulate that only no more than a certain percentage of the total output can be applied to the development of technology and innovation.
@@
Now, taking data as an element of the production function, based on Jones and Tonetti (2020), assume an exponential relationship $\eta$ between data $d(t)$ and technology $z(t)$. Introducing the exogenous variable $\theta$, assume that the policy stipulates that only $\theta$ proportions of output can be dataized and put into production as data elements, i.e.:
@@
It is important to note that the exponential transformation rate relationship between data and technology stems from the nonrivalry and cumulative effects of data. As a factor of production, data can be shared indefinitely without loss, and its value grows exponentially with the scope of its use. In the technical production function, the amount of data directly affects the production efficiency of technology through the exponential form:
$z(t)=d(t)^{\eta}$
$\eta=\left(\frac{logz(t)}{logd(t)}\right)$
Hence, $\eta$ reflects the importance of data to technology. As the amount of data increases, the output efficiency of technology increases at an exponential rate, forming a positive cycle: more data→higher technical efficiency→more output→generate more data. For example, the training of autonomous driving techniques requires a large amount of driving data. Assuming that an algorithm uses 1 million kilometers of driving data with 80% accuracy, when the amount of data is increased to 10 million kilometers (10 times), the accuracy may be increased to 95% due to the cumulative effect of non-competitive data (non-linear growth). This improvement is not linearly cumulative, but rather the data is trained by the algorithm to form an exponential optimization, reflecting the exponential transformation relationship between data and technology.
@@ 3 Numerical Simulation and Phase Diagram Analysis
From Figure 1-A: The effect of changes in $\eta$ and $\theta$ on equilibrium investment $k^{*}$ is linear. An increase in the data technology transformation rate $\eta$ promotes investment, while an increase in the proportion of total output dataization $\theta$ depresses investment. Thus, the boosting effect of $\eta$ on the equilibrium capital stock will gradually slow down with an increase in $\theta$. Figure 1-B shows that the impact of changes in the data technology conversion rate $\eta$ and the proportion of total output dataization $\theta$ on the equilibrium consumption $c^{*}$ is non-linear and has a “mixing” effect. $c^{*}$ is affected by $\eta$ at an inflection point, and the inflection point moves backward with increasing $\theta$. This implies that the promotion of consumption by $\eta$ is not persistent, and there is a possibility that the equilibrium consumption level will decrease at higher stages of data technology conversion. We interpret this as follows:
@@
On the one hand, it is clear from Figure 1-A that technological progress can increase the marginal productivity of capital factors thereby driving the equilibrium capital stock, which is consistent with the rules of reality. And when the dataization ratio of total output increases, from an economic perspective, higher $\theta$ implies that more data are produced, and technology significantly improves and raises total factor productivity, at this time, firms are more biased toward data-driven rather than capital factor inputs for technological improvements, and technological improvement contributes to output partially substituting for capital, coupled with a shift in resources from physical capital investment to data production, leading to a decrease in capital required in the steady state. And, in an environment where technology is highly sensitive to data, i.e., the higher the level of $\eta$, the more pronounced is the impact of the equilibrium capital stock by the increase in the proportion of dataization. As can be seen from Figure 1-B, when $\eta$ is small, an increase in $\theta$ does not have a significant effect on equilibrium consumption. In contrast, when $\eta$ is large, an increase in $\theta$ has the effect of promoting equilibrium consumption. Since $\eta$ is the data-technology conversion rate, which represents the maturity of the mechanism for converting data into technology, the larger $\eta$ is, the more mature the conversion mechanism is, the more the same amount of data will be converted into a higher level of technological achievements, thus more efficiently enabling the transmission of initial data from the production side to the demand side of the individual. Consumers also utilize data-forming technologies to assist in decision-making and effective consumption. Therefore, with a mature mechanism of data technology conversion, increasing the proportion of total macroeconomic output dataization can stimulate consumption growth.
@@
On the other hand, Figure 1-B shows that the overall impact of $\eta$ on equilibrium consumption shows an inverted “U”-shaped trend, which implies that transforming big data into technology is not always effective in stimulating consumption growth, i.e., there exists a threshold for the impact of an increase in the data-technology conversion rate on equilibrium consumption, and the size of this threshold is related to The size of this threshold is related to the proportion of total output dataization $\theta$. Specifically, when $\theta$ is larger, in the domain of definition of $\eta$, the range of promoting effect of increasing $\eta$ on $c^{*}$ will be more persistent, and the threshold of the inverted “U”-shaped trend will be more backward, and gradually evolve into “S”-shaped logistic curve. This may be due to the scale of data is the basis for the establishment of the data-technology transformation mechanism (Learning from Big Data). Therefore, the proportion of total output dataization increases, which means that after the total output is dataized, the volume of data that can be used for technological transformation also increases, and the space and resources for “learning” from data will be expanded, the more reliable the data-technology transformation mechanism is, the higher the success rate of data transformation into technology, and the more accurate the prediction of technology for consumer behavior, thus driving the “resilience” of effective consumption growth.
@@
Further examining Figure 1, we also find that both for $k^{*}$ and $c^{*}$, when $\eta$ is at a higher level ($0.6<\eta<1)$, there is a positive correlation between $\theta$ and $\eta$ at the same level of equilibrium (The Same Color Area in the Figure 1). Therefore, when consumption or capital stock is at the same level of steady state, increasing the macroeconomic total output dataization ratio $\theta$ can effectively promote technical progress $\eta$. Specifically as shown in Figure 2. This will be the theoretical basis for our empirical analysis.
@@
Phase Diagram Analysis. Figures 3, 4 further present the dynamic general equilibrium subject to shocks from changes in $\eta$ and $\theta$ in the macro-economy of this paper, using the phase diagram approach:
@@
Figure 3 presents the impact of an increase in the data-technology conversion rate $\eta$ on consumption and capital stock in general dynamic equilibrium for a constant dataization ratio $\theta$ of total macroeconomic output: The black curve represents the case before the positive shock to $\eta$, $E_{1}$ represents the equilibrium point in that case, and $P_{1}$ represents the inflection point in that case; The blue curve represents the case at the beginning of the positive shock to $\eta$, with a small increase in $\eta$, $E_{2}$ represents the equilibrium point in that case, and $P_{2}$ represents the inflection point in that case; The red curve represents the case at the middle to late stage of the positive shock to $\eta$, with a larger increase in $\eta$, $E_{3}$ represents the equilibrium point, and $P_{3}$ represents the inflection point in that case; The three inflection points $P_{1}$,$P_{2}$,$P_{3}$ are covariant since the aggregate output dataization ratio $\theta$ is constant. When $\eta$ increases, equilibrium capital stock will increase linearly, and the phase diagram is shifted to the right. The blue phase diagram indicates that an increase in $\eta$ before the inflection point value will drive equilibrium consumption growth, and the red phase diagram indicates that an increase in $\eta$ after the inflection point value will slightly reduce equilibrium consumption.
@@
As can be seen from Figures 1-B, the impact of increasing dataization ratio $\theta$ on equilibrium consumption is not apparent in the stage of lower $\eta$. Therefore, the impact of increasing macroeconomic total output dataization ratio $\theta$ on consumption and capital stock in the dynamic general equilibrium in Figure 4 is analyzed for the stage of higher $\eta$: The black curve represents the case before the positive shock of theta, $E_{1}$ represents the equilibrium point in that case, and $P_{1}$ represents the inflection point in that case; The red curve represents the case when $\eta$ is higher as $\theta$ increases, $E_{2}$ represents the equilibrium point in that case, and $P_{2}$ represents the inflection point in that case; The two inflection points, $P_{1}$,$P_{2}$, increase sequentially to the right as the dataization ratio of total output $\theta$ continues to increase. As $\theta$ increases, the equilibrium capital stock will be linearly lower, and the phase diagram is shifted left. The red phase diagram indicates that expanding the proportion of dataization at the mature stage of data-technology conversion will be able to effectively stimulate consumption growth.
From figure 3 we can know that when the dataization ratio $\theta$ is fixed, it means that the threshold of $\eta$ to promote $c^{*}$ growth is fixed, and at this time a small increase in $\eta$ will effectively promote $c^{*}$ growth, while once the threshold is exceeded, continuing to raise $\eta$ will not be conducive to $c^{*}$ growth. And regardless of the magnitude of $\eta$ raising, its effect on $k^{*}$ is always promoting. From Figure 4 we can know that when eta is at a higher stage, raising the proportion of dataization $\theta$ will be able to effectively promote the $c^{*}$ growth and reduce $k^{*}$, at the same time, the threshold value of the effect of $\eta$ raising on the growth of $c^{*}$ will also increase because of $\theta$ increasing.